A 78-inch pipe is cut into two pieces. One piece is five times the length of the other. Find the length of the two pieces.
Answer by bucky(2189)

Let's call the length of the short piece L. Since you are told that the long piece is 5 times as long as the short piece, its length is 5 times L or just 5L.
.
The two pieces added together have to be the length of the original uncut pipe, and you are told that the original pipe was 78 inches before it was cut.
.
So we can add the lengths of the two pieces together and set that sum equal to 78 inches as follows:
.
L + 5L = 78
.
Add the two terms on the left side and you have:
.
6L = 78
.
Solve for L by dividing both sides of the equation by 6 and you get:
.
L = 78/6 = 13
.
So L, the length of the short piece, is 13 inches. The long piece is five times as long as the short piece so the long piece is 5 times 13 and this is 65 inches.
.
As a check you can add the 13 inch piece and the 65 inch piece and see that they do add up to the 78 inches that the original pipe was in length.
.
Hope this explanation helps you to understand the problem better.
.